# On energy and its positivity in spacetimes with an expanding flat de Sitter background

### v1

@@ 2. Main conjecture and theorem
Let $(\Omega,g)$ be a compact Riemannian manifold of nonnegative scalar curvature with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$. Then there exists a constant $\lambda_{\max}\geq 0$ only dependent on $(\Omega,g)$ such that, for any $\lambda\in[0,\lambda_{\rm max}]$, the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and if the mean curvature of $\Sigma$ in $(\Omega,g)$ satisfies $H>2\lambda$, then the energy (17) of the initial data set $(\Omega,g,k=\lambda g,\lambda)$ is well defined and nonnegative:
@@
The constant $\lambda_{\max}$ in Corollary 2.4 is defined as $\alpha_{\lambda}$ in (19) by substituting $E_{\rm LY}$ with its Brown–York counterpart $E_{\rm BY}$ everywhere. Consequently, by the rigidity statement of Shi–Tam [ST02] (see Theorem 3.1 below), $\lambda_{\rm max}>0$ whenever $(\Omega,g)$ is not a subset of Euclidean space $\mathbb{R}^{3}$. For specific examples of this corollary, see Section 5.2.
@@ 3. Positivity Results for Quasi-Local Energies with a Minkowski background
We begin by recalling the classical result concerning isometric embeddings of positively curved 2-surfaces. Let $\Sigma$ be a closed surface with a Riemannian metric of positive Gauss curvature. Then there exists an isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$, known as the Weyl embedding, which bounds a compact convex region $\Omega\subset\mathbb{R}^{3}$. Moreover, the embedding is unique up to Euclidean rigid motions [Nir53, Pog52]. In particular, the mean curvature of the isometric embedding is uniquely determined by the intrinsic metric. Using this fact, Brown and York proposed the following quasi-local mass:
@@
Let $\Omega$ be a compact Riemannian 3-manifold with nonempty boundary $\Sigma=\partial\Omega$ of positive Gauss curvature. Let $H$ be the mean curvature of $\Sigma$ in $\Omega$ with respect to the outward normal, and let $H_{0}$ the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Brown–York energy of $\Sigma$ is defined as
@@
Suppose $\Omega$ is a compact Riemannian 3-manifold with non-negative scalar curvature and nonempty boundary $\Sigma=\partial\Omega$ with positive Gauss curvature and mean curvature. Then the Brown–York energy $E_{\rm BY}$ of $\Sigma$ is non-negative and equals zero if and only if $\Omega$ is isometric to a domain in the Euclidean space $\mathbb{R}^{3}$ and $\Sigma$ has only one component.
Liu and Yau [LY03, LY06] generalized this to a Lorentzian setting, incorporating the full spacetime mean curvature vector.
Definition 3.2 (Liu–Yau Energy).
Let $\Sigma$ be an closed spacelike 2-surface in a spacetime $\mathscr{M}$, and suppose $\Sigma$ has positive Gauss curvature. Let $|\vec{H}|$ denote the norm of the spacetime mean curvature vector of $\Sigma$, and let $H_{0}$ be the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Liu–Yau energy of $\Sigma$ is defined as
@@
If we assume that the surface $\Sigma$ is the boundary of a compact initial data set $(\Omega,g,k)$, then the length of the spacetime mean curvature vector $\vec{H}$ can be computed in terms of the initial data set. Indeed, if the mean curvature vector is spacelike, then the Liu–Yau energy can be expressed as
They established positivity under the dominant energy condition (with $\Lambda=0$).
@@
Suppose $(\Omega,g,k)$ is a compact initial data set satisfying the dominant energy condition (with $\Lambda=0$) and has nonempty boundary $\Sigma=\partial\Omega$ of positive Gauss curvature and $H>|{\rm tr}_{\Sigma}k|$. Then the Liu–Yau energy $E_{\mathrm{LY}}$ of $\Sigma$ is non-negative and equals zero if and only if $(\Omega,g,k)$ can be embedded within Minkowski spacetime $\mathbb{R}^{3,1}$ and $\Sigma$ has only one component.
However, Ó Murchadha, Szabados, and Tod [MST04] constructed examples to show that there exist surfaces in Minkowski space satisfying these assumptions for which both the Brown–York and Liu–Yau masses are strictly positive.
@@ 4. Supporting lemmas and proof of main results
In this section, we prove some supporting lemmas and Theorems 2.3 and 2.4. The first lemma provides a sufficient condition on the Gauss curvature of $\Sigma$ that ensures it can be isometrically embedded inside the cosmological horizon $r=1/\lambda$ in $\mathbb{R}^{3}$. The second and third lemmas address the positivity of the energy $E_{\lambda}$. Finally, we present the proofs of the main theorems by combining the lemmas.
@@ 4.1. Isometric embedding inside the cosmological horizon
The next lemma gives a sufficient (albeit not optimal) condition for $\Sigma$ to lie inside the cosmological horizon $r<1/\lambda$ once embedded into $\mathbb{R}^{3}$.
@@
Let $\Sigma$ be a Riemannian two-sphere. For $\lambda>0$, suppose the Gauss curvature $K$ of $\Sigma$ is bounded below by
@@ 4.2. Positivity of energy
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with boundary $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ satisfying the dominant energy condition for $\Lambda=3\lambda^{2}\geq 0$.
@@
Let $(\Omega,g,k,\lambda)$ be a compact initial data set with some boundary component $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}\geq 0$ holds. If $\lambda\in[0,\alpha_{0}]$, where $\alpha_{0}$ is defined by (30), then the energy (15) is well defined and nonnegative:
@@
By definition, we have $\alpha_{\lambda}\leq\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\rm min}}$ and so $\alpha_{\lambda}\leq\alpha_{0}$, where $\alpha_{0}$ is given by (31). Therefore, since $\lambda\leq\alpha_{\lambda}$ by hypothesis, Lemma 4.1 implies that we embed within the cosmological horizon and Lemma 4.2 implies $E_{\lambda}\geq 0$. The rigidity claim is obtained as in Lemma 4.2 except $C$ ranges from $0<C\leq\frac{2\sqrt{6}}{3\pi}$. ∎
@@
We observe that $\alpha_{\lambda}=\alpha(E_{\rm LY})$ is $\lambda$-dependent, while $\lambda_{\max}=\alpha(E_{\rm BY})$ is not. We claim that $\alpha(E_{\rm LY})\geq\alpha(E_{\rm BY})$. For these values of $\lambda$, note that $E_{\rm LY}\geq E_{\rm BY}$. If $\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}$, then the claim follows immediately from the definition (37). Suppose instead that
@@
where we used (38). This establishes the claim. Therefore, if we choose any $\lambda\in[0,\lambda_{\max}]$ where $\lambda_{\max}=\alpha(E_{\rm BY})$, we have $\lambda\leq\alpha_{\lambda}$ and so by Theorem 2.3, the energy $E_{\lambda}$ is nonnegative. Lastly, if $\lambda_{\max}>0$, then equation (37) implies $\alpha(E_{\rm BY})>0$ and so $E_{\rm BY}>0$. In this case, an argument as in (35) shows that $E_{\lambda}>0$ for all $0<\lambda\leq\lambda_{\rm max}$.
@@ 5. Examples
5.1. Schwarzschild–de Sitter
In this section, we review the Schwarzschild–de Sitter spacetime, a $\Lambda$-vacuum solution to the Einstein equations. First, we show how to express the metric in flat expanding coordinates. Then we study the positivity of energy $E_{\lambda}$ along coordinate spheres. Lastly we determine $\alpha_{\lambda}$.
@@ 5.2. Other $\Lambda$-vacuum examples
In the previous section, we applied Corollary 2.4 to the Schwarzschild–de Sitter spacetime, which is one example of a $\Lambda$-vacuum spacetime. In this section we show that there are many $\Lambda$-vacuum spacetimes where Corollary 2.4 applies.
Let $V$ be a closed Riemannian manifold with positive scalar curvature. Fix $p\in V$ and set $M:=V\setminus\{p\}$. The Green’s function at $p$ for the conformal Laplacian on $V$ exists and is strictly positive. Consequently there is a scalar flat metric $g$ on $M$ which is asymptotically flat with $p$ representing the point at infinity, see e.g., [Sch89, LP87]. Since $R_{g}=0$, the initial data set $(M,g,k=\lambda g,\lambda)$ satisfies the $\Lambda$-vacuum Einstein constraint equations with $\Lambda=3\lambda^{2}$. In the asymptotically flat coordinates $x,y,z$, let $\Sigma_{r}$ be the sphere given by $r^{2}=x^{2}+y^{2}+z^{2}$. For $r$ large enough, $\Sigma_{r}$ has positive Gauss and mean curvatures and bounds a compact region $\Omega_{r}$. Let $\lambda_{\rm max}$ be given by Corollary 2.4. Then for all $\lambda\in[0,\lambda_{\rm max}]$, we can choose $r$ large enough so that $E_{\lambda}$ of $(\Omega_{r},g,k=\lambda g,\lambda)$ is well defined and nonnegative.
@@ 6. Discussion and conclusion
In this work, we addressed the problem of defining energy within the framework of an initial data set admitting a positive cosmological constant. In particular, the presence of a cosmological horizon in the background de Sitter spacetime constitutes an obstruction to a global notion of energy, thereby motivating us to use quasi-local definitions instead. Inspired from the Liu–Yau construction in the asymptotically flat setting, we adapted their definition to the flat expanding patch of de Sitter spacetime:
@@
Several open questions naturally arise from this framework. A first one concerns rigidity: by analogy with the classical Positive Energy Theorem, one expects that the vanishing of the quasi-local energy should characterize data isometric to a domain in the flat expanding de Sitter spacetime. Formally, one may conjecture that
@@
So far, we have obtained only a partial rigidity result: in the rigid case, the boundary $\Sigma=\partial\Omega$ must be round, and this occurs precisely at the critical value $\lambda=\alpha_{\lambda}$. Another open problem concerns the dependence of the quasi-local energy on the parameter $\lambda$. As discussed in Theorem 2.3, our result can be interpreted as identifying a range of admissible values of $\lambda$ for a fixed initial data set $(\Omega,g,k)$. It would be interesting to further investigate how the parameter $\lambda$ interacts with the intrinsic and extrinsic geometry of $\Sigma$. A further promising direction is the development of an optimal embedding formulation. In the asymptotically flat case, Wang and Yau [WY09b, WY09a] introduced a variational quasi-local energy obtained by minimizing over admissible isometric embeddings into Minkowski space. Constructing an analogous variational framework in the de Sitter setting would provide a natural and geometrically consistent extension of our approach.

### v2

@@ 3. Main conjecture and theorem
Fix $\lambda\geq 0$. Let $(\Omega,g)$ be a compact Riemannian manifold of nonnegative scalar curvature with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature $K>0$. Then there exists a constant $\lambda_{\max}\geq 0$ only dependent on $(\Omega,g)$ such that, for any $\lambda\in[0,\lambda_{\rm max}]$, the surface $\Sigma$ isometrically embeds inside the cosmological horizon $r<1/\lambda$ in $\mathbb{R}^{3}$, and if the mean curvature of $\Sigma$ in $(\Omega,g)$ satisfies $H>2\lambda$, then the energy (17) of the initial data set $(\Omega,g,k=\lambda g)$ is well defined and nonnegative:
@@
The constant $\lambda_{\max}$ in Corollary 3.4 is defined as $\alpha$ in (19) by substituting $E_{\rm LY}$ with its Brown–York counterpart $E_{\rm BY}$ everywhere. Consequently, by the rigidity statement of Shi–Tam [ST02] (see Theorem 4.1 below), $\lambda_{\rm max}>0$ whenever $(\Omega,g)$ is not a subset of Euclidean space $\mathbb{R}^{3}$. For specific examples of this corollary, see Section 6.2.
@@ 4. Positivity Results for Quasi-Local Energies with a Minkowski background
We begin by recalling the classical result concerning isometric embeddings of positively curved 2-surfaces. Let $\Sigma$ be a closed surface with a Riemannian metric of positive Gauss curvature. Then there exists an isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$, known as the Weyl embedding, which bounds a compact convex region $\Omega_{0}\subset\mathbb{R}^{3}$. Moreover, the embedding is unique up to Euclidean rigid motions [NIR53, POG52]. In particular, the mean curvature of the isometric embedding is uniquely determined by the intrinsic metric. Using this fact, Brown and York proposed the following quasi-local mass:
@@
Let $(\Omega,g)$ be a compact Riemannian 3-manifold with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature. Let $H$ be the mean curvature of $\Sigma$ in $\Omega$ with respect to the outward normal, and let $H_{0}$ the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Brown–York energy of $\Sigma$ is defined as
@@
Suppose $(\Omega,g)$ is a compact Riemannian 3-manifold with non-empty boundary $\Sigma:=\partial\Omega$ of positive Gauss and mean curvatures. If $(\Omega,g)$ has nonnegative scalar curvature, then the Brown–York energy $E_{\rm BY}$ of $\Sigma$ is nonnegative and equals zero if and only if $\Omega$ is isometric to a domain in the Euclidean space $\mathbb{R}^{3}$ and $\Sigma$ is connected.
Liu and Yau [LY03, LY06] (see also Kijowski [KIJ97]) generalized Definition 4.1 to initial data sets.
Definition 4.2.
Let $(\Omega,g,k)$ be a compact initial data set with nonemtpy boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature. Suppose the mean curvature of $\Sigma$ within $\Omega$ satisfies $H>|\text{tr}_{\Sigma}k|$. Let $H_{0}$ be the mean curvature of the isometric embedding of $\Sigma$ into $\mathbb{R}^{3}$. Then the Liu–Yau energy of $\Sigma$ is defined as
If the initial data set $(\Omega,g,k)$ is embedded within a spacetime, then the assumption on the mean curvature appearing Definition 4.2 implies that the mean curvature vector $\vec{H}$ of $\Sigma$ is spacelike. In fact, we can write the Liu–Yau energy as
@@
From this expression, we see that the definition of $E_{\rm LY}$ does not depend on the specific slice $\Omega$ of spacetime which $\Sigma$ lies in.
Liu and Yau established positivity under the dominant energy condition (with $\Lambda=0$).
@@
Suppose $(\Omega,g,k)$ is a compact initial data set satisfying the dominant energy condition with $\Lambda=0$ and has nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature and $H>|{\rm tr}_{\Sigma}k|$. Then the Liu–Yau energy $E_{\mathrm{LY}}$ of $\Sigma$ is non-negative and equals zero only if $(\Omega,g,k)$ can be embedded within Minkowski spacetime $\mathbb{R}^{3,1}$ and $\Sigma$ is connected.
However, Ó Murchadha, Szabados, and Tod [MST04] constructed examples showing that there exist surfaces in Minkowski space satisfying these assumptions for which both the Brown–York and Liu–Yau masses are strictly positive.
@@
We generalize the example of [MST04] to our de Sitter space setting in section 7.1. This suggests an interesting direction of further research: generalize the Wang–Yau notion of quasi-local energy to the setting of de Sitter space.
@@ 5. Supporting lemmas and proof of main results
In this section, we prove some supporting lemmas and Theorems 3.3 and 3.4. The first lemma provides a sufficient condition on the Gauss curvature of $\Sigma$ that ensures it can be isometrically embedded inside the cosmological horizon $r=1/\lambda$ in $\mathbb{R}^{3}$. The second and third lemmas address the positivity of the energy $E_{\lambda}$. Finally, we present the proofs of the main theorems by combining the lemmas.
@@ 5.1. Isometric embedding inside the cosmological horizon
The next lemma gives a sufficient (albeit not optimal) condition for $\Sigma$ to lie inside the cosmological horizon $r<1/\lambda$ once embedded in Euclidean space $\mathbb{R}^{3}$. The remark that appears after Lemma 5.1 provides an example showing that the condition is not optimal. As the proof shows, the lower bound condition is assumed to appeal to Jung’s theorem. It would be interesting if a more precise condition could be formulated.
@@
Fix $\lambda\geq 0$. Let $\Sigma$ be a Riemannian two-sphere, and suppose the Gauss curvature $K$ of $\Sigma$ is positive and bounded below by
@@ 5.2. Positivity of energy
Let $(\Omega,g,k)$ be a compact initial data set with boundary $\Sigma=\partial\Omega$ of positive Gauss curvature $K>0$ satisfying the dominant energy condition for $\Lambda=3\lambda^{2}\geq 0$.
@@
Fix $\lambda\geq 0$. Let $(\Omega,g,k)$ be a compact initial data set with nonempty boundary $\Sigma:=\partial\Omega$ of positive Gauss curvature $K>0$ and $H>|{\rm tr}_{\Sigma}k|$. Assume that the dominant energy condition with respect to $\Lambda=3\lambda^{2}\geq 0$ holds. If $\lambda\in[0,\alpha_{0}]$, where $\alpha_{0}$ is defined by (30), then the energy (15) is well defined and nonnegative:
@@
By definition, we have $\alpha\leq\frac{2\sqrt{6}}{3\pi}\sqrt{K_{\rm min}}$ and so $\alpha\leq\alpha_{0}$, where $\alpha_{0}$ is given by (31). Therefore, since $\lambda\leq\alpha$ by hypothesis, Lemma 5.1 implies that we embed within the cosmological horizon and Lemma 5.2 implies $E_{\lambda}\geq 0$. The rigidity claim is obtained as in Lemma 5.2 except $C$ ranges from $0<C\leq\frac{2\sqrt{6}}{3\pi}$. ∎
@@
We observe that $\alpha(E_{\rm LY})$ [^4] Note that this is what we called $\alpha$ given in (19). is $\lambda$-dependent, while $\lambda_{\max}=\alpha(E_{\rm BY})$ is not. We claim that $\alpha(E_{\rm LY})\geq\alpha(E_{\rm BY})$. For these values of $\lambda$, note that $E_{\rm LY}\geq E_{\rm BY}$. If $\sqrt{K_{\min}}<\frac{E_{\rm LY}}{r_{\Sigma}^{2}}$, then the claim follows immediately from the definition (37). Suppose instead that
@@
where we used (38). This establishes the claim. Therefore, if we choose any $\lambda\in[0,\lambda_{\max}]$ where $\lambda_{\max}=\alpha(E_{\rm BY})$, we have $\lambda\leq\alpha_{\lambda}$ and so by Theorem 3.3, the energy $E_{\lambda}$ is nonnegative. Lastly, if $\lambda_{\max}>0$, then equation (37) implies $\alpha(E_{\rm BY})>0$ and so $E_{\rm BY}>0$. In this case, an argument as in (35) shows that $E_{\lambda}>0$ for all $0<\lambda\leq\lambda_{\rm max}$. ∎
@@ 6. Examples
6.1. Schwarzschild–de Sitter space
In this section, we review the Schwarzschild–de Sitter space, a $\Lambda$-vacuum solution to the Einstein equations. First, we show how to express the metric in expanding coordinates. Then we study the positivity of energy $E_{\lambda}$ along coordinate spheres. Lastly we determine $\alpha_{\lambda}$.
@@ 6.2. Other $\Lambda$-vacuum examples
In the previous section, we applied Corollary 3.4 to the Schwarzschild–de Sitter space, which is one example of a $\Lambda$-vacuum spacetime. In this section we show that there are many $\Lambda$-vacuum spacetimes where Corollary 3.4 applies.
Let $V$ be a closed Riemannian manifold with positive scalar curvature. Fix $p\in V$ and set $M:=V\setminus\{p\}$. The Green’s function at $p$ for the conformal Laplacian on $V$ exists and is strictly positive. Consequently there is a scalar flat metric $g$ on $M$ which is asymptotically flat with $p$ representing the point at infinity, see e.g., [SCH89, LP87]. Since $R_{g}=0$, the initial data set $(M,g,k=\lambda g)$ satisfies the $\Lambda$-vacuum Einstein constraint equations with $\Lambda=3\lambda^{2}$. In the asymptotically flat coordinates $x,y,z$, let $\Sigma_{r}$ be the sphere given by $r^{2}=x^{2}+y^{2}+z^{2}$. For $r$ large enough, $\Sigma_{r}$ has positive Gauss and mean curvatures and bounds a compact region $\Omega_{r}$. For this sufficiently large $r$, let $\lambda_{\rm max}$ be given by Corollary 3.4. Then for all $\lambda\in[0,\lambda_{\rm max}]$, $E_{\lambda}$ of $(\Omega_{r},g,k=\lambda g)$ is well defined and nonnegative.
@@ 7. Discussion and conclusion
7.1. Dependence on the embedding
We recall that the energy $E_{\lambda}$ depends crucially on the choice of isometric embedding used to define the reference term $\vec{H}_{0}$ (cf. (14)), and hence on the choice of foliation of the reference de Sitter space. In our construction, we use the Weyl embedding theorem to embed $\Sigma$ isometrically into $\mathbb{R}^{3}$, which in turn sits as a umbilical hypersurface in de Sitter space. Had we chosen a different foliation, the associated reference embedding would change, and consequently the value of the energy $E_{\lambda}$ would, in general, be different. A similar phenomenon already appears in the Liu–Yau quasi-local energy, as clearly illustrated by the examples of O’Murchadha, Szabados, and Tod [OST03]. They considered spacelike $2$-surfaces lying on null cones in flat Minkowski space and showed that, despite the fact that the ambient spacetime is globally flat, the Liu–Yau energy of such surfaces can be strictly positive. These examples demonstrate that the value of the quasi-local energy is sensitive to the chosen isometric embedding, rather than being determined solely by the intrinsic geometry of the surface. The same mechanism applies in our setting, as we now describe.
We follow the construction in [CdP21]. Let $\Psi\colon\Sigma\to\mathcal{L}_{\mathscr{M}_{0}}$ be an immersion, where $\Sigma$ is a surface as in Theorem 3.3 and $\mathcal{L}_{\mathscr{M}_{0}}$ denotes the light cone of a point in de Sitter space with cosmological constant $\Lambda=3\lambda^{2}$. By [CdP21, Corollary 4.2], the Gauss curvature $K$ of $\Sigma$ is given by
$K=\frac{1}{4}\lvert\vec{H}_{\Psi}\rvert^{2}+\lambda^{2},$ (50)
where $\vec{H}_{\Psi}$ denotes the mean curvature vector of the image $\Psi(\Sigma)$.
On the other hand, $\Sigma$ can also be isometrically embedded into $\mathbb{R}^{3}$ via the Weyl embedding. Combining (50) with the Gauss equation (see (34)), we obtain
$H_{0}^{2}-4\lambda^{2}-\lvert\vec{H}_{\Psi}\rvert^{2}=2\lvert\mathring{A_{0}}\rvert^{2}\geq 0,$ (51)
[…]
These considerations suggest that a further promising direction is the development of an optimal embedding formulation. In the case of Minkowski space, Wang and Yau [WY09b, WY09a] introduced a variational quasi-local energy defined by minimizing over admissible isometric embeddings into Minkowski space. Constructing an analogous variational framework in the de Sitter setting would provide a natural and geometrically consistent extension of our approach.
7.2. Conclusion
In this work, we addressed the problem of defining energy within the framework of an expanding initial data set with a positive cosmological constant. In particular, the presence of a cosmological horizon in the background de Sitter space constitutes an obstruction to a global notion of energy, thereby motivating us to use quasi-local definitions instead. Inspired from the Liu–Yau construction in the asymptotically flat setting, we adapted their definition to the flat-expanding patch of de Sitter space:
@@ 7.2. Conclusion
Several open questions naturally arise from this framework. A first one concerns rigidity: by analogy with Theorem 4.2, one expects that the vanishing of the quasi-local energy should should imply that the data is isometric to a domain in the flat-expanding patch of de Sitter space. Formally, one may conjecture that
@@
So far, we have obtained only a partial rigidity result: in the rigid case, the boundary $\Sigma=\partial\Omega$ must be round, and this occurs precisely at the critical value $\lambda=\alpha_{\lambda}$. It is an open question whether this rigidity can be upgraded to the full rigidity conjecture stated above.
Another open problem concerns the dependence of the quasi-local energy on the parameter $\lambda$. As discussed in Theorem 3.3, our result can be interpreted as identifying a range of admissible values of $\lambda$ for a fixed initial data set $(\Omega,g,k)$. It would be interesting to further investigate how the parameter $\lambda$ interacts with the intrinsic and extrinsic geometry of $\Sigma$.
Declarations
Conflict of interest
The authors state that there is no Conflict of interest. No data was collected or analysed as part of this project.
Open Access
This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this licence, visit http://creativecommons.org/licenses/by/4.0/.